Two towns A and B ate 42 km apart. A person starts from A,walks at 4 km/h towards B and meets another person coming from B towards A after 6 hours. Find the rate at which the second person is walking, if both persons start walking at the same time.
Answer by htmentor(1343)

Let s = the speed of person B
Since they are walking towards each other, their relative speed is 4 + s.
At this relative speed, they cover a distance of 42 km in 6 hrs.
Therefore 4 + s = 42/6 = 7
Thus s = 7 - 4 = 3 km/h